Tell if the nitrogen balance would be positive or negative if the person takes 95 g of protein in a 24-hr period, and has a urinary urea nitrogen output of 11.6 mg/mL and a 24-hr urine volume of 1200 ml.
Answer #1

Urinary urea nitrogen(UUN) output = 11.6 mg/ml

So,UUN in 1200 ml = 1200/11.6

=103.4 mg

=0.1034 gm

Nitrogen balance = Nitrogen intake - Nitrogen loss

=(Total protein intake in grams / 6.25) - (UUN + 4 grams)

{6.25 = protein per gm of nitrogen &4 gm =nitrogen lost each day}

= 95 / 6.25 -(0.1034 +4 )

= 15.2 - 4.1034

= 11.09 gm

So,the nitrogen balance is positive that is intake is greater than loss.